5. Find a solution u(x, t) of the following problem utt 0 u(0, t) — и(2, t) — 0 2 sin 3T и(а, 0) — 0, и (х, 0) — sin Tz _
Answer #1

Gkven tt phw,h guen predeu i Cltt uhere o S2 boundary aud fal The given Conditons as ulat) ut)-0 0)0, e(,0) = ShIT-2S3 Euatioe Na from uaton,ue have (Coataut Contistet Slutiont 9alioD tathg kp uaten he have IN Kニ0 where DE dot fhear differesid Gaffet Guaion wth Contaat Aendliary Guetion ta giveu by otipti cohere o , B-p ComplinustinyHeute メ= Cf.+PI メ=GspxtGSithpa→6 Azdin fm , we have aT 2pT t (Dap ahere Ds dt difeveucthalyton wth aa Catfhcdeuet gteu bmtpEGasptpt) CF lso hesight laud idet qationg herfre pasrtiudar fileg heule T CP.+PI T- Gos(pt Bpltto or TGlos pt)Ga(ipt) € Nowfom (to), ehave unt)Gnp(pt)+G;f&©¢U® Sfapo Go 78hap SihT, n l,2, 3, --. E 1uLu de n=1,2,3, - - have SroSatien we m,2,3, 2 Po2 snoof m 2,3-- C3 O ation we haue Puty G ult) GC S ( S m2,3, - na12,3 Jaw hypiniplespeyportien, mat genal oluhion is jienbyBs snua)Bm¥) 2 +& sth (3)hlH. +0Sf CuT o Sn ( shcua)tosh( +osh() )25( +osh ( Euoting the Corespordg tesue, howe B-, Be-2 audhaue we utt) bshSE+&,5l) S/2Tt sit1 31Tt 2- (t - ult)ot () sk t0+0 -0t t04 s /a 3T totot 37T heule utrt) )ShCTE) a sduthonfu